wath is the anser to 3/x+1=2/x
Answer by checkley71(8403)

3/X+1=2/X THEN (3+X)/X=2/X OR CROSS MULTIPLYING WE GET 2X=3X+X~2 OR
X~2+X=0 OR X(X+1)=0 OR X=-1
PROOF 3/-1+1=2/-1 OR -3+1=-2 OR -2=-2.